4x + y = 16, 2x + 3y = 24
I have tried mul. the 1st equation by -3. Then adding the 2 together. I have tried mul. 2nd equation by -2, then adding the 2 together. I have tried y = 16 - 4x, then tried to substitute y into the 2nd equation. I can't seem to figure it out.

Answer by Earlsdon(6294)

Solve the system of of equations:
1) 4x+y = 16
2) 2x+3y = 24
Your first step was one in the right direction. Multiply equation 1) by -3 then add the resulting equation to equation 2).
1a) -12x-3y = -48
2)....2x+3y = 24
------------------
-10x = -24 Divide both sides by -10.
x = 2.4 Now substitute this value of x into either one of the two original equations and solve for y.
4x+y = 16 Substitute x = 2.4
4(2.4) + y = 16
9.6 + y = 16 Subtract 9.6 from both sides.
y = 6.4
The solution is (2.4, 6.4)